Four-seventh of a school's faculty are women. Four-fifths of the male faculty members are married, and 9 of the male faculty are unmarried. How many faculty members are there?
Answer by richwmiller(17219)

L+w=f,
4/7f=w,
3/7f=L,
4/5L=m,
1/5L=u,
u=9
faculty f = 105, male L = 45, married m = 36, unmarried male u = 9, women w = 60